# Anosov representations, strongly convex cocompact groups and weak eigenvalue gaps

### v3

@@ 5. Property (U), weak eigenvalue gaps and the uniform gap summation property
In this section, we prove Theorem 1.6 providing conditions under which a representation with a weak uniform gap in eigenvalues is Anosov and we also discuss (strong) property (U), and its relation with the uniform gap summation property.
@@
Let $\Gamma$ be a finitely generated group and fix $|\cdot|_{\Gamma}:\Gamma\rightarrow\mathbb{N}$ a world length function on $\Gamma$. The group $\Gamma$ satisfies property (U) (resp. strong property (U)) if there exists a finite subset $F$ of $\Gamma$ and $C,c>0$ with the following property: for every $\gamma\in\Gamma$ there exists $f\in F$ such that
$\ell_{\Gamma}(f\gamma)\geqslant c|\gamma|_{\Gamma}-C\ \ \big{(}\textup{resp.}\ |f\gamma|_{\infty}\geqslant c|\gamma|_{\Gamma}-C\big{)}.$
@@
Delzant–Guichard–Labourie–Mozes proved in [18] that every word hyperbolic group and every finitely generated group which admits a semisimple quasi-isometric embedding into a reductive Lie group satisfy (strong) property (U). We now prove Theorem 1.7 which implies that virtually torsion free finitely generated groups with non-trivial Floyd boundary [^1] Kassel–Potrie established an analogue of the Abels–Margulis–Soifer lemma [1, Thm. 5.17] simultaneously for a linear representation $\rho:\Gamma\rightarrow\mathsf{GL}_{d}(\mathbb{R})$ of a word hyperbolic group and the abstract group $\Gamma$ equipped with a left invariant word metric (see [31, p. 16]). satisfy strong property (U) (and hence property (U)).
@@
Let $G:[1,\infty)\rightarrow(0,\infty)$ be the function $G(x):=10\sum_{k=\left\lfloor x/2\right\rfloor}^{\infty}f(k)$. Note that $G$ is decreasing and $\underset{x\rightarrow\infty}{\lim}G(x)=0$. By Karlsson’s estimate, see [30, Lem. 1], we have
@@
for every $g,h\in\Gamma$, where $g$ has infinite order. Since $\Lambda(H)$ contains at least $3$ points, by [30, Prop. 5], we may find $f_{1},f_{2}\in H$ non-trivial such that the sets $\{f_{1}^{+},f_{1}^{-}\}$ and $\{f_{2}^{+},f_{2}^{-}\}$ are disjoint. Let us set
$\varepsilon:=\frac{1}{100}\min\big{\{}d_{f}(f_{1}^{+},f_{2}^{+}),d_{f}(f_{1}^{+},f_{2}^{-}),d_{f}\big{(}f_{1}^{-},f_{2}^{+}),d_{f}(f_{1}^{-},f_{2}^{-})\big{\}}$
@@
Claim. 1 Let $F:=\big{\{}f_{1}^{N},f_{2}^{N},e\big{\}}$. Now we claim that for every non-trivial $\gamma\in H$, there exists $g\in F$ such that $d_{f}\big{(}g\gamma^{+},\gamma^{-}\big{)}\geqslant\varepsilon$.
@@
$d_{f}\big{(}\gamma^{n},f_{1}^{-N}\big{)}$ $\geqslant d_{f}\big{(}\gamma^{+},f_{1}^{-}\big{)}-d_{f}\big{(}f_{1}^{-},f_{1}^{-N}\big{)}-d_{f}\big{(}\gamma^{+},\gamma^{n}\big{)}\geqslant 50\varepsilon-G\Big{(}\frac{1}{2}|f_{1}^{N}|_{\Gamma}\Big{)}-G\Big{(}\frac{1}{2}|\gamma^{n}|_{\Gamma}\Big{)}\geqslant 48\varepsilon,$
hence $G\big{(}\big{(}\gamma^{n}\cdot f_{1}^{-N}\big{)}_{e}\big{)}\geq\varepsilon$ for $n\geqslant n_{0}$. By the choice of $M>0$ we have that $\big{(}\gamma^{n}\cdot f_{1}^{-N}\big{)}_{e}\leqslant M$ for $n\geqslant n_{0}$. Then, we choose a sequence $(k_{n})_{n\in\mathbb{N}}$ such that $\big{|}f_{1}^{k_{n}-N}\big{|}_{\Gamma}<\big{|}f_{1}^{k_{n}}\big{|}_{\Gamma}$ for every $n\in\mathbb{N}$. For $n\geqslant n_{0}$ we have
@@
and Claim 1 follows.
Now, let $L:=10\max_{g\in F}|g|_{\Gamma}+2R$. If $\gamma\in H$ and $|\gamma|_{\Gamma}<L$, then we choose $g=e$ and obviously $|\gamma|_{\Gamma}-|\gamma|_{\infty}\leq L$. Suppose that $\gamma\in H$ and $|\gamma|_{\Gamma}\geq L$. We may choose $g\in F$ such that $d_{f}\big{(}(g\gamma g^{-1})^{+},\gamma^{-}\big{)}\geqslant\varepsilon$, where $(g\gamma g^{-1})^{+}=g\gamma^{+}$ in $\partial_{f}\Gamma$. We observe that
@@
$\leqslant G\Big{(}\frac{1}{2}\big{|}\gamma\big{|}_{\Gamma}\Big{)}+G\big{(}(\gamma^{-1}\cdot\gamma^{-1}g^{-1})_{e}\big{)}\leqslant 2G\Big{(}\frac{1}{2}|\gamma|_{\Gamma}-2|g|_{\Gamma}\Big{)}\leqslant\frac{\varepsilon}{50},$
@@
This shows that there is $n_{1}>0$, $G\big{(}\big{(}(g\gamma)^{n}\cdot(g\gamma)^{-1}\big{)}_{e}\big{)}\geqslant\frac{\varepsilon}{3}$ and $\big{(}(g\gamma)^{n}\cdot(g\gamma)^{-1}\big{)}_{e}\leqslant M$ for $n\geqslant n_{1}$. We can find a sequence $(m_{n})_{n\in\mathbb{N}}$ such that
@@
$|g\gamma|_{\Gamma}-|g\gamma|_{\infty}\leqslant 2M\leqslant L.$
This completes the proof of the theorem.∎
@@
Recall that a linear representation $\rho:\Gamma\rightarrow\mathsf{GL}_{d}(\mathbb{R})$ has a weak uniform $i$-gap in eigenvalues if there exists $\varepsilon>0$ such that for every $\gamma\in\Gamma$,
@@
First, we prove the following technical lemma showing that a nilpotent group $\Gamma$ which admits a representation with weak uniform eigenvalue $i$-gap has to be virtually cyclic. [^2] This is not true when $\Gamma$ is assumed to be solvable. The Baumslag–Solitar group $\mathsf{BS}(1,2)$ admits a faithful representation into $\mathsf{GL}_{2}(\mathbb{R})$ with a weak uniform $1$-gap (see [31, Ex. 4.8]).
@@
Let $\Gamma$ be a finitely generated nilpotent group. Suppose that $\rho:\Gamma\rightarrow\mathsf{GL}_{d}(\mathbb{R})$ has a weak uniform $i$-gap in eigenvalues for some $11\leq i\leq d-$ 1. Then $\Gamma$ is virtually cyclic.
@@
Let $G$ be the Zariski closure of $\rho(\Gamma)$ in $\mathsf{GL}_{d}(\mathbb{R})$. We consider the Levi decomposition $G=L\ltimes U$, where $U$ is a connected normal unipotent subgroup of $G$ and $L$ is a reductive Lie group. The projection $\pi\circ\rho:\Gamma\rightarrow L$ is Zariski dense and $\lambda(\pi(\rho(\gamma)))=\lambda(\rho(\gamma))$ for every $\gamma\in\Gamma$. The Lie group $L$ is reductive and $\pi(\rho(\Gamma))$ is solvable, so $L$ has to be virtually abelian since it has finitely many connected components. We may find a finite-index subgroup $H$ of $\Gamma$ such that $\mathfrak{g}_{1}(H)=[H,H]$ is a subgroup of $\textup{ker}(\pi\circ\rho)$. Therefore, for $k\geqslant 1$ we obtain a well defined representation $\rho_{k}:H/\mathfrak{g}_{k}(H)\rightarrow\mathsf{GL}_{d}(\mathbb{R})$ such that $\rho_{k}\circ\pi_{k}=\pi\circ\rho$, where $\pi_{k}:H\rightarrow H/\mathfrak{g}_{k}(H)$ is the quotient map. Note that for every $k\geq 1$ there exists $c_{k}\geq 1$ such that for every $h\in H$,
@@
Since $\lambda(\rho_{k}(h))=\lambda(\rho(h))$ for every $h\in H$, $\rho_{k}$ has a weak uniform $i$-gap in eigenvalues for every $k\geq 1$. We may use induction on $k\in\mathbb{N}$ to see that $H/\mathfrak{g}_{k}(H)$ is virtually cyclic. The group $H/\mathfrak{g}_{1}(H)$ is abelian and satisfies strong property (U), so $\rho_{1}$ is $P_{i}$-Anosov by [31, Prop. 4.12] and $H/\mathfrak{g}_{1}(H)$ has to be virtually cyclic. Now suppose that $H/\mathfrak{g}_{k}(H)$ is virtually cyclic. Note that $\mathfrak{g}_{k}(H)/\gamma_{k+1}(H)$ is a central subgroup of $H/\mathfrak{g}_{k+1}(H)$ with virtually cyclic quotient $H/\mathfrak{g}_{k}(H)$. It follows that $H/\mathfrak{g}_{k+1}(H)$ is virtually abelian. In particular, $H/\mathfrak{g}_{k+1}(H)$ satisfies strong property (U), so $\rho_{k+1}$ is $P_{i}$-Anosov and $H/\mathfrak{g}_{k+1}(H)$ is virtually cyclic. Therefore, $H/\mathfrak{g}_{k}(H)$ has to be virtually cylic for every $k\geq 1$ and $H$ is virtually cyclic since $\mathfrak{g}_{m}(H)=1$ for some $m\geq 1$. ∎
As a corollary of Theorem 1.7, we obtain Corollary 1.8 which shows that a non-virtually nilpotent group $\Gamma$ which admits a representation with the uniform gap summation property satisfies strong Property $U$.
@@
By Proposition 2.4 we may assume that $G=\mathsf{SL}_{d}(\mathbb{R})$ and $\theta=\{\varepsilon_{1}-\varepsilon_{2}\}$. Since $\rho$ satisfies the uniform gap summation property $\textup{ker}(\rho)$ is finite. It suffices to prove that a finite-index subgroup of $\Gamma^{\prime}=\Gamma/\textup{ker}(\rho)$ satisfies strong property (U). By Selberg’s lemma [36] $\Gamma^{\prime}$ is virtually torsion free, so we may assume that $\Gamma$ is torsion free and $\rho$ is faithful. By Lemma 4.8 there exists a continuous $\rho$-equivariant map $\xi_{f}:\partial_{f}\Gamma\rightarrow\mathbb{P}(\mathbb{R}^{d})$ for some Floyd function $f$. We first prove that $\partial_{f}\Gamma$ is not a singleton.
Suppose that $|\partial_{f}\Gamma|=1$. By the definition of the map $\xi_{f}$, the image $\xi_{f}(\partial_{f}\Gamma)$ identifies with the $\tau_{\textup{mod}}$-limit set of $\Gamma$ in $\mathbb{P}(\mathbb{R}^{d})$. Since $\Gamma$ is not virtually nilpotent, we may use [26, Cor. 5.10] to check that $\partial_{f}\Gamma$ contains at least two points. We provide here the following different argument. Since $\partial_{f}\Gamma$ is assumed to be a singleton, up to conjugation, we may assume that $\xi_{f}(\partial_{f}\Gamma)=[e_{1}]$ and find a group homomorphism $a:\Gamma\rightarrow\mathbb{R}^{\ast}$ such that for every $\gamma\in\Gamma$,
$\rho(\gamma)e_{1}=a(\gamma)e_{1}.$
We consider the representation $\hat{\rho}(\gamma)=\frac{1}{a(\gamma)}\rho(\gamma)$. Note that $\hat{\rho}$ satisfies the uniform gap summation property (since $\rho$ does), $\xi_{f}$ is $\hat{\rho}$-equivariant and we can write
$\hat{\rho}(\gamma)=\begin{pmatrix}1&u(\gamma)\\ &\rho_{0}(\gamma)\end{pmatrix}$
@@
$\lim_{n\rightarrow\infty}\frac{\hat{\rho}(g^{n})}{\sigma_{1}(\hat{\rho}(g^{n}))}=k_{\infty}E_{11}k_{\infty}^{\prime}\in\oplus_{i=2}^{d}\mathbb{R}E_{1i}.$
If $\ell_{1}(\hat{\rho}(g))>1$, then $\ell_{1}(\rho_{0}(g))=\ell_{1}(\hat{\rho}(g))$. Let $p_{1}$ and $p_{2}$ be the largest possible dimension of a Jordan block for an eigenvalue of maximum modulus of $\hat{\rho}(g)$ and $\rho_{0}(g)$ respectively. We have
$\sigma_{1}(\hat{\rho}(g^{n}))\asymp n^{p_{1}-1}\ell_{1}(\hat{\rho}(g^{n})),\ \sigma_{1}(\rho_{0}(g^{n}))\asymp n^{p_{2}-1}\ell_{1}(\hat{\rho}(g^{n}))$
@@
$\lim_{n\rightarrow\infty}\frac{1}{n^{p_{1}-1}\ell_{1}(\hat{\rho}(g^{n}))}\sum_{i=0}^{n}i^{p_{2}-1}\ell_{1}(\hat{\rho}(g))^{i}=0.$
Therefore, $\lim_{n}\frac{||u(g^{n})||}{\sigma_{1}(\hat{\rho}(g^{n}))}=0$ which is impossible since $\lim_{n}\frac{\hat{\rho}(g^{n})}{\sigma_{1}(\hat{\rho}(g^{n}))}$ has at least one of its $(1,2),\ldots,(1,d)$ entries non-zero. It follows that $\ell_{1}(\hat{\rho}(g))\leq 1$ and $\ell_{1}(\rho(g))\leq|a(g)|$. Similarly, we obtain $\ell_{d}(\rho(g))^{-1}=\ell_{1}(\rho(g^{-1}))\leq|a(g^{-1})|$. It follows that all the eigenvalues of $\rho(g)$ have modulus equal to $1$. Therefore, by Theorem 2.5, any semisimplification of $\rho$ has compact Zariski closure. Then, by using [2, Thm. 3] and [26, Thm. 10.1], we conclude that $\rho(\Gamma)$ (and hence $\Gamma$) is virtually nilpotent. We have reached a contradiction, therefore, $\xi_{f}$ is non-constant and $\partial_{f}\Gamma$ contains at least two points.
Now we conclude that $\Gamma$ has strong property (U) by showing that $|\partial_{f}\Gamma|\geq 3$.
If $|\partial_{f}\Gamma|=2$, consider the restriction $\rho_{V}:\Gamma\rightarrow\mathsf{GL}(V)$ where $V=\langle\xi_{f}(\partial_{f}\Gamma)\rangle$ and $\dim(V)=2$. We show that all elements of $\rho(\textup{ker}(\rho_{V}))$ have all of their eigenvalues of modulus $1$. For this, since $\xi_{f}(\partial_{f}\Gamma)$ contains two points, up to passing to a finite-index subgroup of $\Gamma$ and conjugating $\rho_{V}$ by an element of $\mathsf{GL}(V)$, we may assume that $\rho_{V}(\Gamma)$ lies in the diagonal subgroup $\mathsf{GL}(V)$. Let $g\in\textup{ker}(\rho_{V})$. We may write $\rho(g^{n})=w_{n}\exp(\mu(g^{n}))w_{n}^{\prime}$ and assume, up to conjugating $\rho$, that, $\lim_{n}w_{n}=w_{\infty},\lim_{n}w_{n}^{\prime}=w_{\infty}^{\prime}$, where $w_{\infty}P_{1}^{+}=P_{1}^{+}$. We see that $\lim_{n}\frac{\rho(g)^{n}}{||\rho(g)^{n}||}=w_{\infty}E_{11}w_{\infty}^{\prime}\in\oplus_{i=1}^{d}\mathbb{R}E_{1i}$ and we may write for $n\in\mathbb{N}$,
$\rho(g^{n})=\begin{pmatrix}I_{2}&\big{(}\sum_{i=0}^{n}A^{i}\big{)}^{t}B\\ &A^{n}\end{pmatrix}$
such that $\lim_{n}\frac{1}{||\rho(g^{n})||}A^{n}$ is the zero matrix. If $A$ has an eigenvalue of modulus greater than $1$, then $\ell_{1}(A)=\ell_{1}(\rho(g))$. By working similarly as in the previous case, we have $\lim_{n}\frac{1}{||\rho(g^{n})||}\sum_{i=0}^{n}||A^{i}||=0$ and $\lim_{n}\frac{1}{||\rho(g^{n})||}\rho(g^{n})$ has all of its $(1,i)$ entries equal to zero, which is absurd. This shows that $\rho(g^{\pm 1})$ has all of its eigenvalues of modulus at most $1$ for $g\in\textup{ker}(\rho_{V})$.
We deduce that $\rho(\textup{ker}(\rho_{V}))$ (and hence $\textup{ker}(\rho_{V})$) is virtually nilpotent and finitely generated. The quotient $\Gamma/\textup{ker}(\rho_{V})$ is abelian, so $\Gamma$ has to be virtually polycyclic. Since $|\partial_{f}\Gamma|>1$, a theorem of Floyd [19, p. 211] implies that $\Gamma$ has two ends, so $\Gamma$ is virtually cyclic. Since $\Gamma$ is assumed not to be virtually nilpotent, this is again a contradiction, hence $\partial_{f}\Gamma$ cannot contain two points.
Finally, it follows that $|\partial_{f}\Gamma|\geq 3$. Therefore, Theorem 1.7 shows that $\Gamma$ satisfies strong property (U). ∎
Proof of Theorem 1.6..
Suppose that (i) holds, i.e $\rho$ is $P_{i}$-Anosov. Then (ii) holds since the Floyd boundary identifies with the Gromov boundary of $\Gamma$. Moreover, by Theorem 2.3 and Proposition 2.6, (iii) and (iv) hold true for any semisimplification $\rho^{ss}$ of the $P_{i}$-Anosov representation $\rho$. Now let us prove the other implications.
We assume that there exists $\varepsilon>0$ such that for every $\gamma\in\Gamma$,
By [31, Prop. 4.12] it is enough to prove that $\Gamma$ satisfies strong property (U).
[…]
The Floyd boundary of $\Gamma^{\prime}=\Gamma/N$ is non-trivial since $\Gamma^{\prime}$ is quasi-isometric to $\Gamma$. Note that the representation $\rho$ induces a faithful representation $\rho^{\prime}:\Gamma^{\prime}\rightarrow\mathsf{GL}_{d}(\mathbb{R})$ which also has a weak uniform $i$-gap in eigenvalues. Selberg’s lemma [36] implies that $\Gamma^{\prime}$ is virtually torsion free so, by Theorem 1.7, $\Gamma^{\prime}$ satisfies strong property (U). We conclude that $\Gamma^{\prime}$ and $\Gamma$ are word hyperbolic and $\rho$ is $P_{i}$-Anosov.

### v4

@@ 5. Property (U), weak eigenvalue gaps and the uniform gap summation property
In this section, we prove Theorem 1.6 providing conditions under which a representation with a weak uniform gap in eigenvalues is Anosov. We also discuss (strong) property (U) and its relation with the uniform gap summation property.
@@
Let $\Gamma$ be a finitely generated group and fix $|\cdot|_{\Gamma}:\Gamma\rightarrow\mathbb{N}$ a word length function on $\Gamma$. The group $\Gamma$ satisfies property (U) (resp. strong property (U)) if there exists a finite subset $F$ of $\Gamma$ and $C,c>0$ with the following property: for every $\gamma\in\Gamma$ there exists $w\in F$ such that
$\ell_{\Gamma}(w\gamma)\geqslant c|\gamma|_{\Gamma}-C\ \ \big(\textup{resp.}\ |w\gamma|_{\infty}\geqslant c|\gamma|_{\Gamma}-C\big).$
@@
Delzant–Guichard–Labourie–Mozes [20] proved that every finitely generated group admitting a semisimple quasi-isometric embedding into a reductive Lie group satisfies (strong) property (U). We now prove Theorem 1.7 which implies that virtually torsion-free finitely generated groups with non-trivial Floyd boundary [^1] Kassel–Potrie established an analogue of the Abels–Margulis–Soifer lemma [1, Thm. 5.17] simultaneously for a linear representation $\rho:\Gamma\rightarrow\mathsf{GL}_{d}(\mathbb{R})$ of a word hyperbolic group and the abstract group $\Gamma$ equipped with a left invariant word metric (see [34, Cor. 1.8]). satisfy strong property (U) (and hence property (U)).
@@
Let $G:[1,\infty)\rightarrow(0,\infty)$ be the function $G(x):=10\sum_{k=\left\lfloor x/2\right\rfloor}^{\infty}f(k)$. Note that $G$ is decreasing and $\underset{x\rightarrow\infty}{\lim}G(x)=0$. By Karlsson’s estimate, see [32, Lem. 1], we have
@@
for every $g,h\in\Gamma$, where $g$ has infinite order. Since $|\Lambda(H)|\geq 3$, by [32, Prop. 5], we may find $f_{1},f_{2}\in H$ infinite order elements such that $\{f_{1}^{+},f_{1}^{-}\}\cap\{f_{2}^{+},f_{2}^{-}\}=\emptyset$. Let us set
$\varepsilon:=10^{-2}\min\big\{d_{f}(f_{1}^{+},f_{2}^{+}),d_{f}(f_{1}^{+},f_{2}^{-}),d_{f}\big(f_{1}^{-},f_{2}^{+}),d_{f}(f_{1}^{-},f_{2}^{-})\big\}>0$
@@
Claim. 1 Let $F:=\big\{f_{1}^{N},f_{2}^{N},e\big\}$. For every non-trivial $\gamma\in H$ there exists $g\in F$ such that $d_{f}\big(g\gamma^{+},\gamma^{-}\big)\geqslant\varepsilon$.
Proof of Claim 1.
@@
$d_{f}\big(\gamma^{n},f_{1}^{-N}\big)$ $\geqslant d_{f}\big(\gamma^{+},f_{1}^{-}\big)-d_{f}\big(f_{1}^{-},f_{1}^{-N}\big)-d_{f}\big(\gamma^{+},\gamma^{n}\big)$
$\geqslant 50\varepsilon-G\Big(\frac{1}{2}|f_{1}^{N}|_{\Gamma}\Big)-G\Big(\frac{1}{2}|\gamma^{n}|_{\Gamma}\Big)\geqslant 48\varepsilon,$
hence $G\big(\big(\gamma^{n}\cdot f_{1}^{-N}\big)_{e}\big)\geq\varepsilon$ for $n\geqslant n_{0}$. By the choice of $M>0$ we have that $\big(\gamma^{n}\cdot f_{1}^{-N}\big)_{e}\leqslant M$ for $n\geqslant n_{0}$. Then, we choose an infinite sequence $(k_{n})_{n\in\mathbb{N}}$ such that $\big|f_{1}^{k_{n}-N}\big|_{\Gamma}<\big|f_{1}^{k_{n}}\big|_{\Gamma}$ for every $n\in\mathbb{N}$. For $n\geqslant n_{0}$ we have
@@
and Claim 1 follows.∎
Now, let $L_{0}:=10\max_{g\in F}|g|_{\Gamma}+2R$. If $\gamma\in H$ and $|\gamma|_{\Gamma}<L_{0}$, then we choose $g=e$ and obviously $|\gamma|_{\Gamma}-|\gamma|_{\infty}\leq L_{0}$. Suppose that $\gamma\in H$ and $|\gamma|_{\Gamma}\geq L_{0}$. We may choose $g\in F$ such that $d_{f}\big((g\gamma g^{-1})^{+},\gamma^{-}\big)\geqslant\varepsilon$, where $(g\gamma g^{-1})^{+}=g\gamma^{+}$ in $\partial_{f}\Gamma$. We observe that
@@
$\leqslant G\Big(\frac{1}{2}|\gamma|_{\Gamma}\Big)+G\big((\gamma^{-1}\cdot\gamma^{-1}g^{-1})_{e}\big)\leqslant 2G\Big(\frac{1}{2}|\gamma|_{\Gamma}-2|g|_{\Gamma}\Big)\leqslant\frac{\varepsilon}{50},$
@@
This shows that there is $n_{1}>0$ with $G\big(\big((g\gamma)^{n}\cdot(g\gamma)^{-1}\big)_{e}\big)\geqslant\frac{\varepsilon}{3}$ and $\big((g\gamma)^{n}\cdot(g\gamma)^{-1}\big)_{e}\leqslant M$ for every $n\geqslant n_{1}$. We can find a sequence $(m_{n})_{n\in\mathbb{N}}$ such that
@@
$|\gamma|_{\Gamma}-|g\gamma|_{\infty}-\big(\max_{g\in F}|g|_{\Gamma}\big)\leq|g\gamma|_{\Gamma}-|g\gamma|_{\infty}\leqslant 2M\leqslant L_{0}.$
In particular, we conclude that $\Gamma$ has strong property (U) and this completes the proof of the theorem.∎
@@
Recall for a matrix $g\in\mathsf{GL}_{d}(\mathbb{R})$, $\ell_{1}(g)\geq\cdots\geq\ell_{d}(g)$ are the moduli of the eigenvalues of $g$. Recall also that a linear representation $\rho:\Gamma\rightarrow\mathsf{GL}_{d}(\mathbb{R})$ has a weak uniform $i$-gap in eigenvalues if there exists $\varepsilon>0$ such that for every $\gamma\in\Gamma$,
@@
First, we prove the following technical lemma showing that a nilpotent group $\Gamma$ which admits a representation with weak uniform eigenvalue $i$-gap has to be virtually cyclic. [^2] This is not true when $\Gamma$ is assumed to be solvable. The Baumslag–Solitar group $\mathsf{BS}(1,2)$ admits a faithful representation into $\mathsf{GL}_{2}(\mathbb{R})$ with a weak uniform $1$-gap (see [33, Ex. 4.8]).
@@
Let $\Gamma$ be a finitely generated nilpotent group. Suppose that $\rho:\Gamma\rightarrow\mathsf{GL}_{d}(\mathbb{R})$ has a weak uniform $i$-gap in eigenvalues for some $1\leq i\leq d-1$. Then $\Gamma$ is virtually cyclic.
@@
Let $G$ be the Zariski closure of $\rho(\Gamma)$ in $\mathsf{GL}_{d}(\mathbb{R})$. We consider the Levi decomposition $G=L\ltimes U$, where $U$ is a connected normal unipotent subgroup of $G$ and $L$ is a reductive Lie group. The projection $\pi\circ\rho:\Gamma\rightarrow L$ is Zariski dense and $\lambda(\pi(\rho(\gamma)))=\lambda(\rho(\gamma))$ for every $\gamma\in\Gamma$. The Lie group $L$ is reductive and $\pi(\rho(\Gamma))$ is solvable, so $L$ has to be virtually abelian since it has finitely many connected components. We may find a finite-index subgroup $H$ of $\Gamma$ such that $\mathfrak{g}_{1}(H)=[H,H]$ is a subgroup of $\textup{ker}(\pi\circ\rho)$. Therefore, for $k\geqslant 1$ we obtain a well-defined representation $\rho_{k}:H/\mathfrak{g}_{k}(H)\rightarrow\mathsf{GL}_{d}(\mathbb{R})$ such that $\rho_{k}\circ\pi_{k}=\pi\circ\rho$, where $\pi_{k}:H\rightarrow H/\mathfrak{g}_{k}(H)$ is the quotient map. Note that for every $k\geq 1$ there exists $c_{k}\geq 1$ such that for every $h\in H$,
@@
Since $\lambda(\rho_{k}(h))=\lambda(\rho(h))$ for every $h\in H$, $\rho_{k}$ has a weak uniform $i$-gap in eigenvalues for every $k\geq 1$. We may use induction on $k\in\mathbb{N}$ to see that $H/\mathfrak{g}_{k}(H)$ is virtually cyclic. The group $H/\mathfrak{g}_{1}(H)$ is abelian and satisfies strong property (U), so $\rho_{1}$ is $P_{i}$-Anosov by [33, Prop. 4.12] and $H/\mathfrak{g}_{1}(H)$ has to be virtually cyclic. Now suppose that $H/\mathfrak{g}_{k}(H)$ is virtually cyclic. Note that $\mathfrak{g}_{k}(H)/\gamma_{k+1}(H)$ is a central subgroup of $H/\mathfrak{g}_{k+1}(H)$ with virtually cyclic quotient $H/\mathfrak{g}_{k}(H)$. It follows that $H/\mathfrak{g}_{k+1}(H)$ is virtually abelian. In particular, $H/\mathfrak{g}_{k+1}(H)$ satisfies strong property (U), so $\rho_{k+1}$ is $P_{i}$-Anosov and $H/\mathfrak{g}_{k+1}(H)$ is virtually cyclic. Therefore, $H/\mathfrak{g}_{k}(H)$ has to be virtually cylic for every $k\geq 1$ and $H$ is virtually cyclic since $\mathfrak{g}_{m}(H)=1$ for some $m\geq 1$. ∎
As a corollary of Theorem 1.7, we obtain Corollary 1.8 which shows that a non-virtually nilpotent group $\Gamma$ which admits a representation with the uniform gap summation property satisfies strong Property (U).
@@
By Proposition 2.4 we may assume that $G=\mathsf{SL}_{d}(\mathbb{R})$ and $\theta=\{\varepsilon_{1}-\varepsilon_{2}\}$. Since $\rho$ satisfies the uniform gap summation property $\textup{ker}(\rho)$ is finite. It suffices to prove that a finite-index subgroup of $\Gamma^{\prime}=\Gamma/\textup{ker}(\rho)$ satisfies strong property (U). By Selberg’s lemma [40], $\Gamma^{\prime}$ is virtually torsion-free, so we may assume that $\Gamma$ is torsion-free and $\rho$ is faithful. By Lemma 4.8 there exists a continuous $\rho$-equivariant map $\xi_{f}:\partial_{f}\Gamma\rightarrow\mathbb{P}(\mathbb{R}^{d})$ for some Floyd function $f$. We first prove that $\partial_{f}\Gamma$ is not a singleton.
Suppose that $|\partial_{f}\Gamma|=1$. By the definition of the map $\xi_{f}$, the image $\xi_{f}(\partial_{f}\Gamma)$ is the $\tau_{\textup{mod}}$-limit set of $\Gamma$ in $\mathbb{P}(\mathbb{R}^{d})$. Since $\Gamma$ is not virtually nilpotent, we may use [28, Cor. 5.10] to check that $\partial_{f}\Gamma$ contains at least two points. We provide here the following different argument that we also use to show that $|\partial_{f}\Gamma|\neq 2$.
Now assume that $\partial_{f}\Gamma$ is a singleton. We shall obtain a contradition. Up to conjugation, we may assume that $\xi_{f}(\partial_{f}\Gamma)=[e_{1}]$ and find a group homomorphism $\alpha:\Gamma\rightarrow\mathbb{R}^{\ast}$ such that for every $\gamma\in\Gamma$,
$\rho(\gamma)e_{1}=\alpha(\gamma)e_{1}.$
We consider the representation $\hat{\rho}(\gamma)=\frac{1}{\alpha(\gamma)}\rho(\gamma)$. Note that $\hat{\rho}$ satisfies the uniform gap summation property (since $\rho$ does), $\xi_{f}$ is $\hat{\rho}-$equivariant and we can write
$\hat{\rho}(\gamma)=\begin{pmatrix}1&u(\gamma)\\ 0&\rho_{0}(\gamma)\end{pmatrix}$
@@
$\lim_{n\rightarrow\infty}\frac{\hat{\rho}(g^{n})}{\sigma_{1}(\hat{\rho}(g^{n}))}=k_{\infty}E_{11}k_{\infty}^{\prime}\in\bigoplus_{i=2}^{d}\mathbb{R}E_{1i}.$
If $\ell_{1}(\hat{\rho}(g))>1$, then $\ell_{1}(\rho_{0}(g))=\ell_{1}(\hat{\rho}(g))$. Let $p_{1}\in\mathbb{N}$ and $p_{2}\in\mathbb{N}$ be the largest possible dimension of a Jordan block for an eigenvalue of maximum modulus of $\hat{\rho}(g)$ and $\rho_{0}(g)$ respectively. A straightforward calculation shows that
$\sigma_{1}(\hat{\rho}(g^{n}))\asymp n^{p_{1}-1}\ell_{1}(\hat{\rho}(g^{n})),\ \sigma_{1}(\rho_{0}(g^{n}))\asymp n^{p_{2}-1}\ell_{1}(\hat{\rho}(g^{n})),\ \ n\rightarrow\infty$
@@
$\lim_{n\rightarrow\infty}\frac{\sum_{i=0}^{n}i^{p_{2}-1}\ell_{1}(\hat{\rho}(g))^{i}}{n^{p_{1}-1}\ell_{1}(\hat{\rho}(g^{n}))}=0.$